# A quick probability-oriented introduction to operator splitting methods

### v1

@@ 7. Some remarks on the relation between ACPs and practice
Generalized examples are [101, 100, 112, 99, 95, 178, 214, 204, 212, 19, 16, 18, 73, 98, 102, 324].
9. Boundary conditions and the phenomenon of order reduction
Now we return to the discussion of the decomposition of a BC.
There are no universal instructions how to proceed here and the following example shows that the corresponding error is unavoidable even in the simplest cases.
@@ 9. Boundary conditions and the phenomenon of order reduction
([188]) Assume that matrices $A_{1},A_{2}\in M_{n}$ commute and consider
@@
Thus the local error is obviously non-zero even for constant $g_{1},g_{2}$ [^34] Or if an attempt to take $g_{1}=g_{2}=g$ is done. However, the method becomes exact if we replace $g_{1},g_{2}$ with
@@
Example 33 shows that one can try to eliminate this source of errors by correcting boundary terms at each step of a splitting scheme. However, (33) contains $\mathrm{e}^{-sA_{2}},$ which means backward integration in time (and solving the initial IVP backward in an abstract setting). Not only this is a problem for implicit methods in general but even such a standard operator as the Laplacian does not permit such integration.
@@
The solution of a one-dimensional heat equation can be written as a series that diverges for negative $t.$ This is a standard simple example of a PDE that cannot be solved backward in time. In particular, MC cannot be used for such backward integration for a diffusion part. However, the heat semigroup is analytic and thus accepts some complex numbers with positive real part.
It appears that the presence of a BC often leads to the phenomenon of order reduction in practice when the rate of the convergence of a splitting scheme is lower that the expected one: e.g. the Strang splitting is generally expected to be of order $2,$ but it happens to be only first order accurate [124].
@@
([124]) An inhomogeneous parabolic ACP in a domain is considered. In the same notation as in Example 33, the corrected Lie-Trotter splitting
@@
is first order accurate. The order of accuracy for the following corrected Strang splitting
@@
admits a bound which is smaller that $2$ unless additional assumptions are satisfied.
Some other references are [176, 117, 118, 276, 116, 275, 9, 10, 115]. In particular, [116] provides a comparison of some methods for the Strang splitting, and [9, 10] study BCs that directly involve the solutions.
@@
Another nontrivial complication is complex BCs that involve time derivative. For the corresponding results, see [87, 9, 10, 217].
@@
This section discusses the general idea behind exponential splitting methods (recall (5.6)) and the Baker-Campbell-Hausdorff (BCH) formula353535also known as the Baker-Campbell-Hausdorff-Dynkin formula in particular.
@@
lives in the Lie group whose Lie algebra is $M_{n}.$ So the local error of the Lie-Trotter and Strang splittings can be written in the terms of Lie brackets ((2.3), (2.4)) and the splitting is exact if the vector fields of the subsystems commute. Note that $\mathrm{e}^{A_{1}}$ and $\mathrm{e}^{A_{2}}$ both live in the same Lie group, which is a crucial moment: the flow of a vector field $A\in C^{\infty}$ defines its own group of diffeomorphisms. Alternatively, we can say $A_{1}$ and $A_{2}$ belong to the same Lie algebra.
This algebraic interpretation goes beyond this finite-dimensional case and is the basis for splitting methods as geometric integrators. References are [169, 40, 252, 188, 296, 297].
Geometric integration can be defined as developing integrators that preserve the underlying geometric properties and structures such as symmetries, solutions living on a special manifold, first integrals etc. An important example is symplectic methods that preserve, in particular, phase volume [^37] They also typically preserve important first integrals and ”almost” preserve energy in practice [296].
@@
(e.g. [171, 40, 297]) Let $H$ be a smooth autonomous Hamiltonian and let $x=(p,q)$ be an element of the corresponding phase space ${\mathbb{R}}^{2m},$ where $p$ are coordinates and $q$ are momenta. The Hamiltonian equations are
@@
We also need the following Baker-Campbell-Hausdorff formula [46, 40, 292, 297] [^39] It is also used in the classical paper [222].. Given elements $a_{1},a_{2}$ of a free Lie algebra the equation in the corresponding Lie group
@@
where all other terms can also be written as combinations of iterated Lie brackets. This expression should be understood as as a formal power series in the corresponding algebra of series [46, 40]. Such abstract interpretation is studied in [15] in the setting of operator splitting methods.
Whether this representation makes sense in a particular setting is actually a nontrivial question. In the case of a finite dimensional Lie algebra $a_{1}$ and $a_{2}$ should be in a neighborhood of $0,$ and the series can diverge even for matrices [41]. Actually, the equation (10.1) is only guaranteed to be always solvable in the algebra of formal polynomials and may not admit the series representation in the terms of commutators of the initial operators etc. [36, 40, 41] [^40] Such pathologies are sometimes ignored in the literature..
@@
and every $B_{i,k},k\geq 1,$ is a linear combination of commutators. Expanding every product of exponentials in powers of $h$ in (10.2) we get
@@
To obtain schemes of the third order and higher some time steps some $b_{i,k}$ and coefficients $a_{i}$ should taken either negative or complex [169, 40] [^41] See also for a discussion of stability [188, 299, 71].. The first condition does not work in the case of diffusion operators (Example 34), while complex time steps may be actually usable here since the semigroup of a diffusion operator can be analytic (in a cone in ${\mathbb{C}}$).
@@
([188]) If $S_{h}$ denotes one step of the Strang splitting, the scheme whose one step equals
@@
The idea to use the BCH formula and the approach of formal power series in the case of PDEs usually fails [40] as the corresponding smooth algebraic structure is not evident or does not exist, [^42] This is related to the notion of hypoellipticity and the Hörmander condition. See also [35, 45]. and the applicability of the BCH formula is not clear. Numerical approximations are usually discretized, as we have already discussed, and thus do not share properties of the real solution. The possible presence of BCs and incompatible domains which the operators of the system and subsystems are defined on are other problems. Still, one can the results formally or as the first steps of a rigorous proof (which usually involves bounds for commutators), or simply use the method to develop new higher order exponential schemes.
The Lie formalism [169, 227, 188, 106] [^43] The Lie formalism can be used to handle the Taylor expansion (e.g. [98]). is typically used then. That is, given an ODE or a IVP
@@
Now any exponential splitting schemes can be treated in the terms of such exponentials. In particular, we can get Examples 30 and 31 since the commutator $[L_{f_{1}},L_{f_{2}}]$ is calculated at $\mathrm{Id}.$ That is, we need $(L_{f_{1}}L_{f_{2}}-L_{f_{2}}L_{f_{1}})\mathrm{Id}$).
@@
A general framework for unbounded generators of $C_{0}-$semigroups is developed in [177]. Other references include [86, 125, 38, 341, 180, 315, 316, 317, 318, 315, 42, 39, 212, 71]. Sources of the physical nature in Section 8 are also often valid here.
Algebraic calculation around the BCH formula and error expansions of Section 8 can be combined (e.g. [212]).
@@
[163] develops higher order schemes by combining simpler operator splitting methods on meshes with different sizes and with variable time steps444444which does not pair well with the BCH formula, for instance for possibly degenerate second order parabolic equations in Sobolev spaces.
@@
The backward error analysis for exponential splitting methods is given in [169]. This universal approach represents a numerical scheme as the flow of some equation and uses a truncated version of this flow to study the global error.
@@
So-called B-series can be used instead of the BCH formalism [169, 40, 15].
@@ 11. The Itô-Taylor expansion, the Kunita expansion and exponential methods for SDEs
Roughly speaking, classical numerical methods for strong approximations of SDEs are typically based on iterating the Itô formula and the Itô-Taylor (Wagner-Platen) expansion (alternatively, the Stratonovich-Taylor expansion) that follows, and weak approximations add parabolic equations, the Girsanov theorem and the Malliavin calculus into the picture. However, the Malliavin calculus and the theory of Lie algebras are extensively involved in the well-established stochastic algebraic theory behind the Stratonovich-Taylor expansion and its truncated versions [222, 320, 255, 223, 224, 136] (also e.g. [245, 270, 339, 134, 274, 338, 199]) [^45] The number of publications devoted to the algebraic interpretation of the Itô-Taylor expansion and the corresponding advances in the Malliavin calculus is huge so all sources listed in this section should merely be seen as starting points or semirandom examples. in general and in obtaining cubature formulas for weak approximations [241, 83, 85, 167, 181] in particular [^46] Cubature formulas and so called Kusuoka approximations are well known to be closed related internally (cf. [223] and [241], for instance)., which provides an appropriate counterpart of the theory in the previous section.
@@
where $A$ is the set of specially defined indexes, $\{c_{\alpha}\}$ are numbers, $\{I_{\alpha}\}$ are iterated Stratonovich integrals w.r.t. $w_{0},\ldots,w_{m}$ and $\{X^{\alpha}\}$ are iterated commutators of $X_{0},\ldots,X_{m}.$ By a classical result of H. Kunita, the series is convergent when the Lie algebra $L$ generated by $X_{0},\ldots,X_{m}$ is nilponent. On the other hand, this representation can formally be understood as the exponential mapping on some non-commutative Lie algebra (say, the algebra of smooth operators). In practice, a truncated version of $\eta$ after removing higher order commutators is considered.
Another approach uses truncated versions of the Stratonovich-Taylor expansion (roughly speaking, this produces so called Kusuoka approximations, in particular). Since it is hard to calculate such iterated integrals efficiently, cubature formulas replace mathematical expectations of such integrals with iterated integrals w.r.t a family of functions of bounded variation. To find cubature formulas explicitly, Lie algebras are typically used, too.
So algebraic calculations are one important component needed to construct such weak approximations. For instance, one can still construct exponential splitting schemes. However, the nature of calculations in general can be rather different unless one indeed combines simpler schemes (such as the Ninomiya-Victoir scheme): though exponential mappings $\mathrm{e}^{y_{n}}$ that correspond to small time cubature formulae or approximate schemes appear, their series are typically truncated so algebraic calculations are often performed in the terms of such finite sums over some symbolic algebras (cf. [15] where such truncations are also studied).
Lower order schemes may not even need the Malliavin calculus or symbolic calculations if a test function is smooth and has bounded derivatives [270].
[255, 274, 320, 136, 256] directly discuss what can be seen as stochastic composition and exponential splitting methods. [133] provides a different perspective in the spirit of the rough path theory.
@@
([320, 270]) The well known Ninomiya-Victoir scheme can actually be seen as an exponential splitting scheme. Recall that for

### v2

@@ 7. Some remarks on the relation between ACPs and practice
General examples are [DesSchatz02Strangs, DesMa04operator, DuMaDes11parareal, DesDuLou07local, Des01convergence, HanOs16high, KochNeuThal13error, JahnLu00error, KochLu08variational, AuKochThal12defect_1, AuHofKochThal15defect_3, AuKochThal15defect_local, CharMeThalZhang16improved, DesDuDuLouMa11adaptive, DesThal13Lie, Thal08high].
9. Boundary conditions, inhomogeneous ACPs and the phenomenon of order reduction
Now we return to the discussion of the decomposition of BC.
In general, BC often leads to the phenomenon of order reduction when the rate of the convergence of a splitting scheme is observed to be lower that expected [HundsVer07Numerical]: e.g. the order of the Strang splitting is strictly lower than $2$ (it may be only first order accurate in practice [FaouOsSchratz15analysis]), while the Lie-Trotter splitting may happen to preserve the first order in the same setting.
Recall that a (semi-)discretized version of a PDE in a domain can be considered as an inhomogeneous ODE where BC is explicitly used to derive the RHS. Thus the following example shows that a naive and straightforward approach leads to unavoidable additional errors even in the simplest cases both in the context of inhomogeneous PDEs (ODEs) and non-trivial BC (as well as the corresponding numerical schemes for such equations).
@@ 9. Boundary conditions, inhomogeneous ACPs and the phenomenon of order reduction
([HundsVer07Numerical]) Assume that matrices $A_{1},A_{2}\in M_{n}$ commute and consider
@@
Thus the local error is obviously non-zero even for constant $g_{1},g_{2}.$ However, the method becomes exact if we replace $g_{1},g_{2}$ with
@@
On the other hand, the theoretical analysis of ACPs often incorporates BC into the definitions of the corresponding domains, so they do not appear in the equation directly. Otherwise a corrective additional inhomogeneous term appears.
@@
([EinOs16overcoming_2]) Consider the Laplacian in a domain $D$ with IC $u_{0}$ and BC $f.$ The solution can be represented as $v+w$ where $w$ is the harmonic extension of $f$ in $D$ and $v$ is the solution of
$\frac{\partial v}{\partial t}=\Delta v+g(v,w),$
$v(0)=u_{0}-w_{0}\quad\mathrm{in\ }D,$
$v=0\quad\mathrm{on\ }\partial D,$
with the correction $g$ [^38] See [FaouOsSchratz15analysis] for a discussion of the situation when $w$ is problematic to find.. However, $\frac{\partial v}{\partial t}=g$ do not necessarily satisfy the original BC and may thus be incompatible with the Laplacian (in the sense of domains), which is a problem for the theoretical error analysis.
Remark 48.
Note that this point of view can be somewhat different from most typical approaches based on the weak formulation and Galerkin-Faedo approximations where BC is often imposed onto ODEs for basis coefficients or is included via regularization. For a quick summary of the relation between inhomogeneous PDEs, non-trivial BC, the weak formulation of a PDE, Galerkin-Faedo appproximations and other common numerical methods, see e.g. references in [PeetPeet24treatment]. See also [LuSiWahl92semigroup] for the semigroup approach.
Example 33 shows that one can try to eliminate this source of errors by correcting boundary terms at each step of a splitting scheme. However, (33) contains $\mathrm{e}^{-sA_{2}},$ which means backward integration in time (solving the initial IVP backward in an abstract setting). This is not only a problem for implicit methods in general but even such a standard operator as the Laplacian does not permit such integration. Obviously, the same remarks apply to the decomposition of the inhomogeneous part of a PDE in general, even without BC.
@@
The solution of a one-dimensional heat equation can be written as a series that diverges for negative $t.$ This is a standard simple example of a PDE that cannot be solved backward in time. However, the heat semigroup is analytic and thus accepts some complex numbers with positive real part.
Example 36.
([FaouOsSchratz15analysis]) Dimension splitting for an inhomogeneous two-dimensional parabolic PDE393939in the divergence form, which is typical for such analytical results in a domain is considered. For an inhomogeneous ACP and in the same notation as in Example 33, the corrected Lie-Trotter splitting
@@
is first order accurate under suitable regularity assumptions on $g.$ The order of accuracy for the corrected Strang splitting
@@
is guaranteed to be less than $2$ unless additional assumptions on the smoothness of $g$ and $u(0)$ are imposed.
Some other references are [HanOs09dimension, EinOs15overcoming_1, EinOs16overcoming_2, OsSchratz13stability, EinOs18comparison, OsSchratz12error, AltZi23secondArx, AltKoZi23bulk, EinMoOs18efficient, HanKraOs12second]. In particular, [EinOs18comparison] provides a comparison of some methods for the Strang splitting, and [AltZi23secondArx, AltKoZi23bulk] study BC that directly involve the solutions. In particular, [OsSchratz13stability] develops general results on stability of splitting schemes in the terms of smoothing properties of analytic semigroups involved (that is, in the terms of fractional powers of the corresponding generators); domains of various operators that appear in this context should also be suitably compatible [^40] Cf. the discussion of the control for domains in the proof of the Trotter-Karo formula in Section 4.. Typically, the BC and inhomogeneous part of the equation should be sufficiently smooth. A problem of BC with time derivatives is considered in [CsoEhrFas21operator, AltZi23secondArx, AltKoZi23bulk, KoLu17numerical].
@@
A very naive – and obvious – explanation of some results here is as follows: to obtain higher order stability results, one needs to expand a solution (given e.g. via the variation-of-constants formula or via (8.3)) and estimate the remainder, which yields assumptions on the regularity of the solution. This produces requirements for BC and IC, and it is exactly where fractional powers of closed operators appear. See e.g. [OsSchratz13stability] for a detailed example. Probabilists also use such techniques in the theory of SPDEs and in the theory of subordinated Feller and sub-Markovian semigroups [But20Method, Ja02pseudoVol2, Ja02pseudoVol3] [^41] Subordination is a random time change by a subordinator., for example. In particular, see [But20Method, Ja02pseudoVol2] and other related references in Section 4 for applications of the Trotter-Kato formula in the latter context.
Remark 50.
Some detailed numerical splitting schemes can be found e.g. in [EinMoOs18efficient].
@@
This section discusses the general idea behind exponential splitting methods (recall (5.6)) and the Baker-Campbell-Hausdorff (BCH) formula424242also known as the Baker-Campbell-Hausdorff-Dynkin formula in particular.
@@
lives in the Lie group whose Lie algebra is $M_{n}.$ So the local error of the Lie-Trotter and Strang splittings can be written in the terms of Lie brackets ((2.3), (2.4)) and the splitting is exact if the vector fields of the subsystems commute. Note that $\mathrm{e}^{A_{1}}$ and $\mathrm{e}^{A_{2}}$ both live in the same Lie group, which is a crucial moment: the flow of a vector field $A\in C^{\infty}$ defines its own group of diffeomorphisms in the general case of non-constant coefficients. Alternatively, we can say $A_{1}$ and $A_{2}$ belong to the same Lie algebra.
This algebraic interpretation goes beyond this finite-dimensional case and is the basis for splitting methods as geometric integrators. References are [HaiLu10geometric, BlaCa16concise, McLachQuis02splitting, HundsVer07Numerical, Sanz97geometric, SanzCal94numerical].
Geometric integration can be defined as developing integrators that preserve some underlying geometric properties and structures such as symmetries, solutions living on a special manifold, first integrals etc. An important example is symplectic methods that preserve, in particular, phase volume [^44] They also typically preserve important first integrals and ”almost” preserve energy in practice [Sanz97geometric].
@@
(e.g. [Hall13quantum, BlaCa16concise, SanzCal94numerical]) Let $H$ be a smooth autonomous Hamiltonian and let $x=(p,q)$ be an element of the corresponding phase space ${\mathbb{R}}^{2m},$ where $p$ are coordinates and $q$ are momenta. The Hamiltonian equations are
@@
We also need the following Baker-Campbell-Hausdorff formula [BonFul12theorem, BlaCa16concise, Ross02introduction, SanzCal94numerical] [^46] It is also used in the classical paper [Ku80representation].. Given elements $a_{1},a_{2}$ of a free Lie algebra the equation in the corresponding Lie group
@@
where all other terms can also be written as combinations of iterated Lie brackets. This expression should be understood as a formal power series in the corresponding algebra of series [BonFul12theorem, BlaCa16concise]. Such an abstract interpretation is studied in [AuHerKochThal17BCH_formula] in the setting of operator splitting methods.
Whether this representation makes sense in a particular setting is actually a non-trivial question. In the case of a finite dimensional Lie algebra $a_{1}$ and $a_{2}$ should be in a neighborhood of $0,$ and the series can diverge even for matrices [BlaCa04convergence]. Actually, the equation (10.1) is only guaranteed to be always solvable in the algebra of formal polynomials and may not admit the series representation in the terms of commutators of the initial operators etc [BiaBonMa20Baker, BlaCa16concise, BlaCa04convergence] [^47] Such pathologies are sometimes ignored in the literature..
@@
and every $B_{i,k},k\geq 1,$ is a linear combination of commutators. Expanding every product of exponentials in powers of $h$ in (10.2) gives
@@
To obtain schemes of the third and higher order some time steps $b_{i,k}$ and coefficients $a_{i}$ should be either negative or complex [HaiLu10geometric, BlaCa16concise] [^48] See also for a discussion of stability [HundsVer07Numerical, Schatz02numerical, CasCharDesVil09splitting].. The first condition does not work in the case of diffusion operators (Example 35), while complex time steps may be actually usable here since the semigroup of a diffusion operator can be analytic (in a cone in ${\mathbb{C}}$).
@@
([HundsVer07Numerical]) If $S_{h}$ denotes one step of the Strang splitting, the scheme whose one step equals
@@
The idea to use the BCH formula and the approach of formal power series in the case of PDEs usually fails [BlaCa16concise] as the corresponding smooth algebraic structure is not evident or does not exist, [^49] This is related to the notion of hypoellipticity and the Hörmander condition. See also [BiaBon14convergence, Bon10ode_version]. and the applicability of the BCH formula is not clear. Numerical approximations are usually discretized, as we have already discussed, and thus do not share geometrical properties of the real solution. The possible presence of BC and incompatible domains which the operators of the system and subsystems are defined on are other problems. Still, one can apply results formally or as the first steps of a rigorous proof (which usually involves bounds for commutators), or simply use the method described above to find new higher order exponential schemes.
The Lie formalism [HaiLu10geometric, LanVer99analysis, HundsVer07Numerical, DiFaHaZla01commutativity] [^50] The Lie formalism can be used to handle the Taylor expansion (e.g. [DesDuDuLouMa11adaptive]). is typically used then. That is, given an ODE or a IVP
@@
Now any exponential splitting schemes can be treated in the terms of such exponentials. In particular, we obtain Examples 30 and 31 since the commutator $[L_{f_{1}},L_{f_{2}}]$ is calculated at $\mathrm{Id},$ so we use $(L_{f_{1}}L_{f_{2}}-L_{f_{2}}L_{f_{1}})\mathrm{Id}$.
@@
A general framework for unbounded generators of $C_{0}-$semigroups is developed in [HanOs09exponential]. Other references include [CsoFaHa05weighted, FaGei07iterative, BlaCalCaSanz21symmetrically, Yo90construction, HaSu05finding, Su91general, Su92general, Su95hybrid, Su93improved, Su91general, BlaCaMu08splitting, BlaCaCharMu13optimized, KochLu08variational, CasCharDesVil09splitting]. Sources of the physical nature in Section 8 are also often valid here.
Algebraic calculation around the BCH formula and error expansions of Section 8 can be combined (e.g. [KochLu08variational]).
@@
[GyKry05accelerated] develops higher order schemes by combining simpler operator splitting methods on meshes with different sizes and with variable time steps515151which does not pair well with the BCH formula, for instance for possibly degenerate second order parabolic equations in Sobolev spaces.
@@
The backward error analysis for exponential splitting methods is given in [HaiLu10geometric]. This universal approach represents a numerical scheme as the flow of some equation and uses a truncated version of this flow to study the global error.
@@
So-called B-series can be used instead of the BCH formalism [HaiLu10geometric, BlaCa16concise, AuHerKochThal17BCH_formula].
@@ 11. The Itô-Taylor expansion, the Kunita expansion and exponential methods for SDEs
Roughly speaking, classical numerical methods for strong approximations of SDEs are typically based on iterating the Itô formula and the corresponding Itô-Taylor (Wagner-Platen) expansion (alternatively, the Stratonovich-Taylor expansion), and weak approximations add parabolic equations, the Girsanov theorem and the Malliavin calculus into the picture. However, the Malliavin calculus and the theory of Lie algebras are extensively involved in the well-established stochastic algebraic theory behind the Stratonovich-Taylor expansion and its truncated versions [Ku80representation, TaKo09operator, Mi01Lie, Ku01approximation, Ku04approximation, Fu06sixth] (also e.g. [MaNohSil20algebraic, NoVic08weak, Ya18weak, FourLasLeLions99applications, OshiTeichVe12new, Ya17higher, IguYa21operator]) [^52] The number of publications devoted to the algebraic interpretation of the Itô-Taylor expansion and the corresponding advances in the Malliavin calculus is huge so all sources listed in this section should merely be seen as starting points or semirandom examples. in general and in obtaining cubature formulas for weak approximations [LyonsVic04cubature, CriMaNee13cubature, CriOr13Kusuoka, GyurLyons11efficient, HaTa22MC_construction] in particular, which provides an appropriate counterpart of the theory in the previous section.
@@
where $A$ is the set of special indexes, $\{c_{\alpha}\}$ are numbers, $\{I_{\alpha}\}$ are iterated Stratonovich integrals w.r.t. $w_{0},\ldots,w_{m}$ and $\{X^{\alpha}\}$ are iterated commutators of $X_{0},\ldots,X_{m}.$ By a classical result of H. Kunita, the series is convergent when the Lie algebra $L$ generated by $X_{0},\ldots,X_{m}$ is nilponent. On the other hand, this representation can formally be understood as the exponential mapping on some non-commutative Lie algebra (say, the algebra of smooth operators). In practice, higher order commutators are discarded and a truncated version of $\eta$ is considered.
Another approach uses truncated versions of the Stratonovich-Taylor expansion (roughly speaking, this produces so called Kusuoka approximations, in particular [^53] Cubature formulas and so called Kusuoka approximations are well known to be closed related internally (cf. [Ku01approximation] and [LyonsVic04cubature], for instance).). Since it is hard to calculate such iterated integrals efficiently, cubature formulas replace mathematical expectations of such integrals with iterated integrals w.r.t a family of functions of bounded variation. To find cubature formulas explicitly, Lie algebras are typically used, too.
So algebraic calculations are one important component needed to construct such weak approximations. For instance, one can still construct exponential splitting schemes. However, the nature of calculations in general can be rather different unless one indeed combines simpler schemes (such as the Ninomiya-Victoir scheme): though exponential mappings that correspond to small time cubature formulae or approximate schemes appear, the series are typically truncated so algebraic calculations are often performed in the terms of finite sums over some symbolic algebras (cf. [AuHerKochThal17BCH_formula] where such truncations are also studied).
Lower order schemes may not even need the Malliavin calculus or symbolic calculations if a test function is smooth and has bounded derivatives [NoVic08weak].
[Mi01Lie, OshiTeichVe12new, TaKo09operator, Fu06sixth, Mi00numerical] directly discuss what can be seen as stochastic composition and exponential splitting methods. [FosReisStrange23highArx] provides a different perspective in the spirit of the rough path theory.
@@
([TaKo09operator, NoVic08weak]) The well known Ninomiya-Victoir scheme can actually be seen as an exponential splitting scheme. Recall that for